# TF-Mamba: A Time-Frequency Network for Sound Source Localization

### v1

@@ Abstract
Sound source localization (SSL) determines the position of sound sources using multi-channel audio data. It is commonly used to improve speech enhancement and separation. Extracting spatial features is crucial for SSL, especially in challenging acoustic environments. Previous studies performed well based on long short-term memory models. Recently, a novel scalable SSM referred to as Mamba demonstrated notable performance across various sequence-based modalities, including audio and speech. This study introduces the Mamba for SSL tasks. We consider the Mamba-based model to analyze spatial features from speech signals by fusing both time and frequency features, and we develop an SSL system called TF-Mamba. This system integrates time and frequency fusion, with Bidirectional Mamba managing both time-wise and frequency-wise processing. We conduct the experiments on the simulated dataset and the LOCATA dataset. Experiments show that TF-Mamba significantly outperforms other advanced methods on simulated and real-world data.
Index Terms:
@@
I Introduction
Sound source localization (SSL) [1] aims to automatically identify the origins of sound by analyzing signals captured through a microphone array. The core functionality of SSL systems is to calculate the angles at which sounds reach the microphones. This spatial information is essential for many speech-based applications, including speech separation [2, 3, 4], speech recognition [5], and speech enhancement [6]. Accurate sound localization can assist these applications to improve their performance. As a result, developing more advanced SSL techniques has become a key focus in research, aiming for greater robustness and adaptability in real-world.
@@ I Introduction
In recent years, deep learning methods for localization have gained more attention [10, 11, 12] than traditional methods. These methods approach the task as either feature/location regression or location classification. Common architectures for moving SSL include convolutional neural networks (CNNs) [13, 14] and recurrent neural networks (RNNs) [15, 16]. CNNs extract local spatial features, while RNNs capture long-term temporal context. Deep learning models can take input at the signal level (e.g., time-domain signals, spectrograms) or feature level (e.g., inter-channel phase difference, generalized cross-correlation (GCC) [17], spatial spectrum). Motivated by speech enhancement, the FN-SSL [18] framework builds on previous work, recognizes that the direct-path inter-channel phase difference can be estimated in both narrow-band (frequencies) and full-band (time-frames). It uses dedicated LSTM layers to process each band and show advanced performance.
@@
While LSTM-based models like FN-SSL have shown promising results, a new neural state space model (SSM)-based architecture called Mamba [19, 20] has emerged and achieves superior performance to state-of-the-art models across various tasks [21, 22]. In speech processing, there have been attempts to replace transformers and RNNs with Mamba for tasks like speech separation and speech enhancement [23, 24, 25]. Mamba demonstrates a stronger ability to model extremely long-range dependencies compared to LSTM and is notable for its efficient use of computational resources.
In this work, we introduce Mamba to SSL by proposing a novel architecture referred to as TF-Mamba. It is built on the robust FN-SSL framework, which uses LSTM for time and frequency feature fusion. By replacing the LSTM blocks in FN-SSL with bidirectional Mamba (BiMamba) blocks, TF-Mamba aims to improve audio sequence context modeling while maintaining linear complexity with sequence length. We consider both simulated and LOCATA datasets for our studies to evaluate the effectiveness of proposed TF-Mamba for real-world scenarios. To the best of our knowledge, this is the first attempt to use a state-space-based model for SSL.
@@
Figure 1: Architecture of the proposed TF-Mamba network. Each TF-Mamba block includes a temporal Mamba (T-Mamba) and frequency Mamba (F-Mamba) layer, with skip connections to prevent information loss. Bidirectional Mamba (BiMamba) layers capture both past and future information. “S” denotes the SiLU activation and “Linear” indicates linear projection.
II Related Works
II-A Mamba for Speech
@@
Mamba has demonstrated transformer-level performance across various sequence-based modalities, including audio and speech, which are naturally sequential in waveform or spectrogram forms. Early works applied Mamba to single tasks such as speech enhancement [26, 25], speech separation [23, 24], and audio detection and classification [27, 28]. Additionally, some studies explored self-supervised audio transformers trained with masked spectrogram modeling [29]. The most comprehensive studies to date examine Mamba’s applications in speech enhancement, recognition, synthesis, understanding, and summarization. However, effective model design using SSMs for deep SSL remains unexplored.
II-B Deep Sound Source Localization
@@
III Proposed TF-Mamba
III-A Mamba
The structured SSMs [19, 20] efficiently handles long-dependent sequences with low computation and memory needs, serving as a substitute for transformers or RNNs. Mamba [34] enhances SSM by introducing an input-dependent selection mechanism for efficient information filtering and a hardware-aware algorithm that scales linearly with sequence length, enabling faster computation. Mamba’s architecture, combining SSM blocks with linear layers, is simpler but achieves state-of-the-art performance across various long-sequence tasks, offering significant computational efficiency during both training and inference.
The core of Mamba is a linear selective SSM. In the equation below, the variables $h_{t}$, $x_{t}$, and $y_{t}$ represent the hidden state, input, and output at time $t$, respectively. The matrices $A$, $B$, and $C$ are learnable parameters corresponding to the state transition, input, and output processes.
$h_{t}=Ah_{t-1}+Bx_{t},\quad y_{t}=Ch_{t}$ (1)
Due to its linear nature, the entire output sequence $y$ of length $\mathcal{L}$ can be expressed as a convolution between the input sequence $x$ of the same length and a kernel $\mathcal{K}$.
$\quad y=x\ast\mathcal{K},{\text{~{}where~{}}}\mathcal{K}=(CB,AB,\ldots,CA^{\mathcal{L}-1}B)$ (2)
Since the matrices $A_{t}$, $B_{t}$, and $C_{t}$ depend on the input $x_{t}$ (i.e., they are selective), this convolution cannot be computed directly but is instead solved using a parallel scan algorithm. A unidirectional Mamba is an SSM positioned between gated linear layers. For most speech-related tasks, bidirectional modeling is preferred because it captures information from both past and future sequences. Therefore, we use the bidirectional Mamba (BiMamba) proposed in [25], where two SSMs and causal convolutions run in parallel: one processes the original sequence, and the other processes the reversed sequence. The outputs from both SSMs are averaged to incorporate information from both directions.
III-B Network Architecture
The structure of the proposed TF-Mamba network is illustrated in Fig. 1. The input to the network consists of the real and imaginary parts of the STFT coefficients, resulting in an input channel count that is twice the number of microphones. The model has three modules as follows:
III-B1 Feature Encoder
@@
III-B2 Time-Frequency Mamba
The Time-Frequency Mamba (TF-Mamba) block consists of one temporal Mamba layer and one frequency Mamba layer. Fig. 1 illustrates a network with a TF-Mamba block, with additional blocks easily added by repeating the block multiple ($N$) times.
@@ III-B2 Time-Frequency Mamba
Mamba is useful because it efficiently models long-range dependencies while maintaining computational efficiency. The ability of Mamba to handle both temporal and frequency-specific features in a structured manner makes it ideal for capturing complex relationships in audio data, such as the time-varying nature of localization cues, which are crucial for accurate SSL. However, the temporal Mamba and frequency Mamba layers are designed to focus on their respective types of information. Therefore the pieces of information from the frequency Mamba layer can be lost after passing through a temporal Mamba layer, and vice versa. To prevent this information loss, we add skip connections. As shown in Fig. 1, the output of each temporal Mamba layer is added to the input of the following temporal Mamba layer. Similarly, the output of each frequency Mamba layer is added to the input of the next frequency Mamba layer. Further, each Mamba layer inside BiMamba also has the skip connection.
III-B3 Output Decoder
@@
IV Experiment Setting
IV-A Datasets
@@
IV-B Implementation details and metrics
@@
TABLE I: Ablation studies of TF-Mamba on simulated data. “Expand” denotes the expand factor for the Mamba layer.
Components Params[M] ACC($15^{\circ}$) [%] ACC($10^{\circ}$) [%] MAE []
@@
IV-C Reference baselines
We compared the following five moving SSL methods: Cross3D [30], SELDnet [32], SE-Resnet [31], SALSA-Lite [33], and FN-SSL [18]. It is noted that the SELDnet, SALSA-Lite, and SE-ResNet are models designed for joint sound event detection and localization; hence, we only utilize their localization branches for comparison. All these methods are reproduced and trained on the same dataset as our proposed method. We adjust the time pooling kernel size in each method to achieve the same output frame rate across all models for consistency.
TABLE II: Performance comparison in ACC and MAE on simulated data.
Methods Params Clean -10 dB
[M] ACC($15^{\circ}$) ACC($10^{\circ}$) MAE[] ACC($15^{\circ}$) ACC($10^{\circ}$) MAE[]
SELDnet 0.8 96.6 93.8 3.5 52.8 50.2 16.7
FN-SSL 2.1 98.3 95.9 2.8 71.6 68.4 8.9
Cross3D 5.6 84.8 81.3 6.5 38.8 35.4 26.3
SE-ResNet 10.2 96.7 95.1 3.2 67.7 64.6 10.8
SALSA-Lite 14.0 96.8 95.6 2.9 69.1 65.7 10.3
TF-Mamba 1.8 98.9 96.9 2.5 74.6 72.5 8.2
TABLE III: Performance comparison in ACC and MAE on LOCATA dataset.
Methods Params[M] ACC($15^{\circ}$) ACC($10^{\circ}$) MAE []
@@ IV-C Reference baselines
TF-Mamba 1.8 97.2 94.3 3.2
V Results and Analysis
V-A Ablation: Finding the best configuration for TF-Mamba
@@
The ablation study results for TF-Mamba, shown in Table I, highlight the importance of different components in the network. “Block $\times$ $N$” refers to stacking $N$ number of TF-Mamba blocks. As the number of blocks increases from 3 to 5, accuracy ($10^{\circ}$ error tolerances) improves from 92.6% to 96.9%, showing that deeper networks help the model understand more complex localization information. The “w/o BiMamba” variant, which replaces the bidirectional mamba with unidirectional, reduces accuracy to 93.6% ($10^{\circ}$) and 97.1% ($15^{\circ}$). This shows that processing both past and future information is important for achieving good performance. The “w/o Skip” variant, which excludes skip connections, also leads to a small drop in accuracy and a higher error. This confirms that skip connections are helpful in keeping information flowing through the network. We use Block $\times$ 5 version with BiMamba and skip connections in the following comparison experiments.
V-B Results on simulated data
@@
We report the results in two acoustic conditions of testing in Table II which are “Clean” and “SNR=-10 dB”. The results demonstrate that TF-Mamba outperforms other methods in clean and noisy environments, highlighting its robustness in handling noise and reverberation. With the highest accuracy in $10^{\circ}$ (96.9%), $15^{\circ}$ (98.9%) and lowest MAE (2.5) in clean conditions, TF-Mamba outperforms FN-SSL. Although FN-SSL also analyzes direct-path features, TF-Mamba captures time and frequency dependencies more effectively, leading to better accuracy and robustness in challenging environments. In contrast, methods like SALSA-Lite, which processes noisy IPD with magnitude spectrum, and Cross3D, which uses the SRP-PHAT spatial spectrum, show significant performance drops under noise, underscoring the advantage of TF-Mamba.
V-C Results on real-world data
@@
The results on the LOCATA dataset are presented in Table III, where the models trained on simulated data are directly tested. We observe that TF-Mamba achieves the best performance, with the highest accuracy at both 15° (97.2%) and 10° (94.3%) error tolerances, and the lowest MAE (3.2). This dataset’s relatively good acoustic conditions, which are almost noise-free and have an RT60 of 0.55s, allow all models to perform well. However, TF-Mamba’s superior accuracy and lower error indicate its ability to better capture fine localization details compared to the other methods. Notably, methods like SE-ResNet and SALSA-Lite, which are more complex in terms of parameters, show higher MAE, while simpler models like SELDnet struggle with accuracy and MAE. TF-Mamba’s approach also avoids the over-smoothing problem seen in other models during sudden sound source turns, maintaining strong localization performance in real-world conditions.
VI Conclusion

### v2

@@ Abstract
Sound source localization (SSL) determines the position of sound sources using multi-channel audio data. It is commonly used to improve speech enhancement and separation. Extracting spatial features is crucial for SSL, especially in challenging acoustic environments. Recently, a novel structure referred to as Mamba demonstrated notable performance across various sequence-based modalities. This study introduces the Mamba for SSL tasks. We consider the Mamba-based model to analyze spatial features from speech signals by fusing both time and frequency features, and we develop an SSL system called TF-Mamba. This system integrates time and frequency fusion, with Bidirectional Mamba managing both time-wise and frequency-wise processing. We conduct the experiments on the simulated and real datasets. Experiments show that TF-Mamba significantly outperforms other advanced methods. The code will be publicly released in due course.
keywords:
@@
1 Introduction
Sound source localization (SSL) [1] aims to automatically identify the origins of sound by analyzing signals captured through a microphone array. The core functionality of SSL systems is to calculate the angles at which sounds reach the microphones. This spatial information is essential for many speech-based applications, including speech separation [2, 3, 4], speech recognition [5], and speech enhancement [6]. Accurate sound localization can assist these applications to improve their performance. As a result, developing more advanced SSL techniques has become a key focus in research, aiming for greater robustness and adaptability in the real world.
@@ 1 Introduction
In recent years, deep learning methods for localization have gained more attention [10, 11, 12] than traditional methods. These methods approach the task as either feature/location regression or location classification. Common architectures for moving SSL include convolutional neural networks (CNNs) [13, 14] and recurrent neural networks (RNNs) [15, 16]. CNNs extract local spatial features, while RNNs capture long-term temporal context. Deep learning models can take input at the signal level (e.g., time-domain signals, spectrograms) or feature level (e.g., inter-channel phase difference, spatial spectrum). Inspired by speech enhancement, the recent FN-SSL [17] framework builds on previous research by effectively estimating the localization of the sound source in both the narrow-band (frequency) and full-band (timeframe) domains. Long short-term memory (LSTM) layers are used to process each band, demonstrating the state-of-the-art performance.
@@
Although RNN models like FN-SSL have shown promising results, a new architecture based on the neural state space model (SSM) called Mamba [18, 19] has emerged and achieves superior performance to state-of-the-art models in various tasks [20, 21, 22]. In speech processing, there have been attempts to replace transformers and RNNs with Mamba for tasks such as speech separation and speech enhancement [23, 24, 25]. As [26] state, Mamba outperforms RNNs in modeling very long-range dependencies and is recognized for its efficient use of computational resources which is suitable for SSL applications.
In this work, we introduce Mamba to SSL by developing a novel architecture called TF-Mamba. Our goal is to enhance SSL by utilizing Mamba’s capability to efficiently model long-range dependencies. Additionally, we apply Mamba to both temporal and frequency dimensions, creating a dual-dimension approach that improves feature extraction and better captures spatial and localization cues compared to baseline models. By utilizing bidirectional Mamba (BiMamba) blocks, TF-Mamba enhances audio sequence context modeling while maintaining computational complexity. We evaluate TF-Mamba using both simulated data and the LOCATA dataset, demonstrating its effectiveness in real-world scenarios. To our knowledge, this is the first study to apply a state-space-based model to SSL, showcasing the novelty and significance of our approach.
@@
Figure 1: Architecture of (a) the proposed TF-Mamba network and (b) Bidirectional Mamba (BiMamba) layer. Each TF-Mamba block includes a temporal Mamba (T-BiMamba) and frequency Mamba (F-BiMamba) layer, with skip connections to prevent information loss. “S” denotes the SiLU activation and “Linear” indicates linear projection.
2 Related Works
2.1 Mamba for Speech
@@
Mamba [26] has demonstrated transformer-level performance across various sequence-based modalities, including audio and speech, which are naturally sequential in waveform or spectrogram forms. Early works applied Mamba to single tasks such as speech enhancement [25, 27], speech separation [23], and audio detection and classification [28, 29]. Additionally, some studies explored self-supervised audio transformers trained with masked spectrogram modeling. The most comprehensive studies to date examine Mamba’s applications in speech enhancement, recognition, synthesis, understanding, and summarization. However, effective model design using SSMs for deep SSL remains unexplored.
2.2 Deep Sound Source Localization
@@
3 Proposed TF-Mamba
3.1 BiMamba
The structured SSMs [18, 19] efficiently handle long-dependent sequences with low computation and memory needs, serving as a substitute for transformers or RNNs. Mamba [26] enhances SSM by introducing an input-dependent selection mechanism for efficient information filtering and a hardware-aware algorithm that scales linearly with sequence length, enabling faster computation. Mamba’s architecture, combining SSM blocks with linear layers, is simpler but achieves state-of-the-art performance across various long-sequence tasks, offering significant computational efficiency during both training and inference.
The core of Mamba is a linear selective SSM. The SSM transforms an input $x_{t}$ into an output $y_{t}$ using a hidden state $h_{t}$ at each step $t$ in the sequence. It includes three learnable matrices: $A$ for state transitions, $B$ for inputs, and $C$ for outputs. A discretization process is applied to continuous-time SSMs for the integration into practical deep neural architectures. We convert the continuous matrices $A$ and $B$ into discrete forms, $\tilde{A}$ and $\tilde{B}$. This step allows the model to work with discrete-time signals. After this, the SSM follows the equation shown below:
$h_{t}=\tilde{A}h_{t-1}+\tilde{B}x_{t},\quad y_{t}=Ch_{t}$ (1)
Because of its linear structure, the full output sequence $y$ of length $\mathcal{L}$ can be written as a convolution between the input sequence $x$ and a kernel $\mathcal{K}$. Each element of $\mathcal{K}$ shows how an input at one time step affects future outputs through the state transitions. This kernel summarizes the model’s temporal behavior and allows efficient processing of long sequences. This makes structured state-space models useful for SSL.
$\quad y=x\ast\mathcal{K},{\text{~{}where~{}}}\mathcal{K}=(C\tilde{B},C\tilde{A}\tilde{B},\ldots,C\tilde{A}^{\mathcal{L}-1}\tilde{B})$ (2)
However, the matrices in the model depend on the input $x_{t}$, making them input-selective. Because of this, the convolution cannot be computed directly. Instead, it is solved using a parallel scan algorithm. In the Mamba model, the SSM is placed between two gated linear layers. These added components improve performance by introducing selective features, but the model remains unidirectional, like a traditional RNN. While unidirectionality is acceptable in large language models that use autoregressive training, it becomes a problem in non-autoregressive speech models. For the SSL task, models must capture both past and future context. Therefore, a non-causal module, such as attention, is needed. Finding an effective way to add this capability is an important challenge.
We use the bidirectional Mamba (BiMamba) proposed in [25, 22], where two SSMs and causal convolutions run in parallel: one processes the original sequence, and the other processes the reversed sequence. The outputs from both SSMs are averaged to incorporate information from both directions.
3.2 Network Architecture
The structure of the proposed TF-Mamba network is illustrated in Figure 1 (a). The input to the network consists of the real and imaginary parts of the STFT coefficients, resulting in an input channel count that is twice the number of microphones. The model has three modules as follows:
3.2.1 Feature Encoder
@@
3.2.2 Time-Frequency Mamba
The Time-Frequency Mamba (TF-Mamba) block consists of one temporal Mamba layer and one frequency Mamba layer. Figure 1 (a) illustrates a network with a TF-Mamba block, with additional blocks easily added by repeating the block multiple ($N$) times. Figure 1 (b) illustrates the BiMamba layer.
@@ 3.2.2 Time-Frequency Mamba
Mamba is useful because it efficiently models long-range dependencies while maintaining computational efficiency. The ability of Mamba to handle both temporal and frequency-specific features in a structured manner makes it ideal for capturing complex relationships in audio data, such as the time-varying nature of localization cues, which are crucial for accurate SSL. However, the temporal Mamba and frequency Mamba layers are designed to focus on their respective types of information. Therefore the pieces of information from the frequency Mamba layer can be lost after passing through a temporal Mamba layer, and vice versa. To prevent this information loss, we add skip connections. As shown in Figure 1, the output of each temporal Mamba layer is added to the input of the following temporal Mamba layer. Similarly, the output of each frequency Mamba layer is added to the input of the next frequency Mamba layer. Further, each Mamba layer inside BiMamba also has the skip connection.
3.2.3 Output Decoder
@@
4 Experiment Setting
4.1 Datasets
@@
4.2 Implementation details and metrics
@@
4.3 Reference baselines
We compared the following five moving SSL methods: Cross3D [30], SELDnet [32], SE-Resnet [31], SALSA-Lite [33], and FN-SSL [17]. It is noted that the SELDnet, SALSA-Lite, and SE-ResNet are models designed for joint sound event detection and localization; hence, we only utilize their localization branches for comparison. All these methods are reproduced and trained on the same dataset as our proposed method. We adjust the time pooling kernel size in each method to achieve the same output frame rate across all models.
Table 1: Ablation studies of TF-Mamba on simulated data. “Expand” denotes the expand factor for the Mamba layer.
Components Params[M] $\downarrow$ ACC($15^{\circ}$) $\uparrow$ ACC($10^{\circ}$) $\uparrow$ MAE [] $\downarrow$
@@ 4.3 Reference baselines
Table 2: Performance comparison in ACC and MAE on simulated data. We test the models under the clean set and with -10 dB noise. The highlight is our proposed TF-Mamba model.
Methods Params[M] $\downarrow$ Clean ACC($15^{\circ}$) $\uparrow$ ACC($10^{\circ}$) $\uparrow$ MAE[] $\downarrow$
SELDnet 0.8 96.6 93.8 3.5
FN-SSL 2.1 98.3 95.9 2.8
Cross3D 5.6 84.8 81.3 6.5
SE-ResNet 10.2 96.7 95.1 3.2
SALSA-Lite 14.0 96.8 95.6 2.9
[…]
SELDnet 0.8 52.8 50.2 16.7
FN-SSL 2.1 71.6 68.4 8.9
Cross3D 5.6 38.8 35.4 26.3
SE-ResNet 10.2 67.7 64.6 10.8
SALSA-Lite 14.0 69.1 65.7 10.3
[…]
Table 3: Performance comparison in ACC and MAE on LOCATA dataset. The highlight is our proposed TF-Mamba model.
Methods Params[M] $\downarrow$ ACC($15^{\circ}$) $\uparrow$ ACC($10^{\circ}$) $\uparrow$ MAE [] $\downarrow$
@@
[…]
5 Results and Analysis
5.1 Ablation: Configuration for TF-Mamba
@@
The ablation study results for TF-Mamba, shown in Table 1, highlight the importance of different components in the network. “Block $\times$ $N$” refers to stacking $N$ number of TF-Mamba blocks. As the number of blocks increases from 3 to 5, accuracy ($10^{\circ}$ error tolerances) improves from 92.6% to 96.9%, showing that deeper networks help the model understand more complex localization information. The “w/o BiMamba” variant, which replaces the bidirectional mamba with unidirectional, reduces accuracy to 93.6% ($10^{\circ}$) and 97.1% ($15^{\circ}$). This shows that processing both past and future information is important for achieving good performance. The “w/o Skip” variant, which excludes skip connections, also leads to a small drop in accuracy and a higher error. This confirms that skip connections are helpful in keeping information flowing through the network. We use Block $\times$ 5 version with BiMamba and skip connections in the following comparison experiments.
5.2 Results on simulated data
@@
We report the results in two acoustic conditions of testing in Table 4.3 which are “Clean” and “SNR=-10 dB”. The results demonstrate that TF-Mamba outperforms other methods in clean and noisy environments, highlighting its robustness in handling noise and reverberation. With the highest accuracy in $10^{\circ}$ (96.9%), $15^{\circ}$ (98.9%) and lowest MAE (2.5) in clean conditions, TF-Mamba outperforms FN-SSL by more effectively capturing long-range time and frequency dependencies, resulting in higher accuracy and greater robustness in challenging environments. This improvement demonstrates our contribution by validating the effectiveness of dual-dimension Mamba modeling. In contrast, methods like SALSA-Lite, which processes noisy IPD with magnitude spectrum, and Cross3D, which uses the SRP-PHAT spatial spectrum, show significant performance drops under noise, underscoring the advantage of TF-Mamba.
5.3 Results on real-world data
@@
The results on the LOCATA dataset are presented in Table 4.3, where the models trained on simulated data are directly tested. We observe that TF-Mamba achieves the best performance, with the highest accuracy at both 15° (97.2%) and 10° (94.3%) error tolerances, and the lowest MAE (3.2). This dataset’s relatively good acoustic conditions, which are almost noise-free and have an RT60 of 0.55s, allow all models to perform well. However, TF-Mamba’s superior accuracy and lower error indicate its ability to better capture fine localization details compared to the other methods. Notably, methods like SE-ResNet and SALSA-Lite, which are more complex in terms of parameters, show higher MAE, while simpler models like SELDnet struggle with accuracy and MAE. TF-Mamba’s approach also avoids the over-smoothing problem seen in other models during sudden sound source turns, maintaining strong localization performance in real-world conditions.
6 Conclusion